OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 36 when x = 9??

jimthompson5910 (jim_thompson5910):

y = kx 36 = k*9 Solve for k

jimthompson5910 (jim_thompson5910):

oh wait, i misread y = k/x 36 = k/9 there fixed